OpenStudy (anonymous):

Fan and Medal Solve. u^2 - 22u = 9

OpenStudy (texaschic101):

have you tried the quadratic formula ?

OpenStudy (anonymous):

he said yes

OpenStudy (imstuck):

You would have to use the quadratic formula to solve this. Or complete the square, but if you don't know how to complete the square, trying to teach it in this forum is almost impossible! Do you know the quadratic formula?\[\frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\]and fill in your a, b, and c.\[\frac{ -(-22)\pm \sqrt{(-22)^{2}-4(1)(-9)} }{ 2(1) }\]That would all simplify to \[\frac{ 22\pm \sqrt{484-36} }{ 2 }\]which, simplified even further is \[\frac{ 22\pm \sqrt{448} }{ 2 }\]Now 448 is not a perfect square, so you need to break it up into its components.\[\sqrt{448}=\sqrt{64\times7}\]That has a perfect square of 8 in it (the square root of 64 is 8), so we now have\[\frac{ 22\pm8\sqrt{7} }{ 2 }\]Now divide everything by the denominator of 2 to get \[11\pm4\sqrt{7}\]So your two solutions are \[11+4\sqrt{7} \]and\[11-4\sqrt{7}\]See that?

OpenStudy (texaschic101):

check your math... 22 +- sq rt (484 + 36) / 2

OpenStudy (anonymous):

Is it 11+sq rt 130